# Late-time X-ray Observations of the Jetted Tidal Disruption Event AT2022cmc: The Relativistic Jet Shuts Off

### v1

@@ 3.5 Constraining the Mass Accretion Rate and the Total Accreted Mass
While the precise nature of relativistic jet production in TDEs is not well understood, such jets are canonically expected to form when the accretion rate onto the black hole exceeds the Eddington rate (Giannios & Metzger, 2011). The low occurrence rate of jetted TDEs however suggests that super-Eddington accretion, which is expected at early times for disruptions of solar mass stars by black holes with $M_{\rm BH}\lesssim 10^{8}\ \rm M_{\odot}$ (De Colle et al., 2012), is not a sufficient condition for powering relativistic jets, and that additional parameters such as black hole spin or disk-jet alignment must play a role (Stone & Loeb, 2012; Tchekhovskoy et al., 2014; Franchini et al., 2015; Curd & Narayan, 2019; Zanazzi & Lai, 2019; Teboul & Metzger, 2023). Nevertheless, the rapid drop in flux observed in the X-ray light curves of jetted TDEs on timescales commensurate with a transition to sub-Eddington accretion as seen in numerical simulations supports a connection between super-Eddington accretion and jet production (De Colle et al., 2012).
@@ 3.6 Revival of the Jet
As the fallback accretion rate continues to decline following $\dot{M}\propto t^{-2.1}$ and eventually reaches a few percent of Eddington, the disk is expected to transition to a radiatively inefficient advection-dominated accretion flow (Maccarone, 2003), analagous to the “low/hard” state in X-ray binaries (Fender et al., 2004; Russell et al., 2011). At this point, the jet is expected to turn back on. Based on the Eddington accretion rate derived in Section 3.5, we estimate that the mass accretion rate will reach $2\%$ Eddington at $t_{\rm obs}\approx 3-5$ years (or $t_{\rm obs}\approx 5-9$ years if instead the mass accretion rate more closely follows $\dot{M}\propto t^{-5/3}$) with an associated X-ray flux of $F_{X}\approx 1-5\times 10^{-16}\ \rm erg\ cm^{-2}\ s^{-1}$. As this is beyond the detection limits of current X-ray facilities with reasonable exposure times, we conclude that continued X-ray monitoring of AT2022cmc is unlikely to detect a revival of the jetted emission, unless the transition to a “low/hard” state occurs at a larger fraction of the Eddington accretion rate. On the other hand, such a revival may be accompanied by a rapid rise in radio flux from the forward shock interaction of the jet with the ambient medium. Indeed, late-time radio rebrightening has been observed in a large fraction ($\sim 40\%$) of optically-selected TDEs (Horesh et al., 2021; Cendes et al., 2023), although the underlying mechanism governing the rebrightening, as well as whether the mechanism is the same for all TDEs, remains unclear.
@@ Appendix A Constraining the Jet Shut-off Time in Sw J2058+05
The scientific results reported in this article are based on observations made by the Chandra X-ray Observatory under program GO 24700310, PI: Eftekhari. T. E. is supported by NASA through the NASA Hubble Fellowship grant HST-HF2-51504.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. A. T. acknowledges support by NSF AST-2009884, AST-2107839, AST-1815304, AST-1911080, OAC-2031997, AST-2206471 and NASA 80NSSC22K0031 and 80NSSC21K1746 grants.

### v2

@@ 3.5 Constraining the Mass Accretion Rate and the Total Accreted Mass
While the precise nature of relativistic jet production in TDEs is not well understood, such jets are canonically expected to form when the accretion rate onto the black hole exceeds the Eddington rate (Giannios & Metzger, 2011). The low occurrence rate of jetted TDEs however suggests that super-Eddington accretion, which is expected at early times for disruptions of solar mass stars by black holes with $M_{\rm BH}\lesssim 10^{8}\ \rm M_{\odot}$ (De Colle et al., 2012), is not a sufficient condition for powering relativistic jets, and that additional parameters such as black hole spin or disk-jet alignment must play a role (Stone & Loeb, 2012; Tchekhovskoy et al., 2014; Franchini et al., 2016; Curd & Narayan, 2019; Zanazzi & Lai, 2019; Teboul & Metzger, 2023). Nevertheless, the rapid drop in flux observed in the X-ray light curves of jetted TDEs on timescales commensurate with a transition to sub-Eddington accretion as seen in numerical simulations supports a connection between super-Eddington accretion and jet production (De Colle et al., 2012).
@@ 3.6 Revival of the Jet
As the fallback accretion rate continues to decline following $\dot{M}\propto t^{-2.1}$ and eventually reaches a few percent of Eddington, the disk is expected to transition to a radiatively inefficient advection-dominated accretion flow (Maccarone, 2003), analagous to the “low/hard” state in X-ray binaries (Fender et al., 2004; Russell et al., 2011). At this point, the jet is expected to turn back on. Based on the Eddington accretion rate derived in Section 3.5, we estimate that the mass accretion rate will reach $2\%$ Eddington at $t_{\rm obs}\approx 3-5$ years (or $t_{\rm obs}\approx 5-9$ years if instead the mass accretion rate more closely follows $\dot{M}\propto t^{-5/3}$) with an associated X-ray flux of $F_{X}\approx 1-5\times 10^{-16}\ \rm erg\ cm^{-2}\ s^{-1}$. As this is beyond the detection limits of current X-ray facilities with reasonable exposure times, we conclude that continued X-ray monitoring of AT2022cmc is unlikely to detect a revival of the jetted emission, unless the transition to a “low/hard” state occurs at a larger fraction of the Eddington accretion rate. On the other hand, such a revival may be accompanied by a rapid rise in radio flux from the forward shock interaction of the jet with the ambient medium. Indeed, late-time radio rebrightening has been observed in a large fraction ($\sim 40\%$) of optically-selected TDEs (Horesh et al., 2021b; Cendes et al., 2023), although the underlying mechanism governing the rebrightening, as well as whether the mechanism is the same for all TDEs, remains unclear.
While jets have not been observed to turn back on for the existing sample of relativistic TDEs, late-time Chandra observations of Sw J1644+57 have detected faint X-rays consistent with emission from the expanding forward shock (Zauderer et al., 2013; Levan et al., 2016; Eftekhari et al., 2018). X-ray observations of AT2022cmc at late times, coupled with an extrapolation of the radio SED into the X-ray regime, will probe such emission and enable constraints on the synchrotron cooling frequency. Finally, we note that the detection of quiescent X-ray emission from a pre-existing low-luminosity AGN is unlikely, given the distance to AT2022cmc and expected luminosities for low-luminosity AGN of $L_{X}\lesssim 10^{42}\rm erg\ s^{-1}$ (Brusa et al., 2007).
@@ Appendix A Constraining the Jet Shut-off Time in Sw J2058+05
The scientific results reported in this article are based on observations made by the Chandra X-ray Observatory under program GO 24700310, PI: Eftekhari. T. E. is supported by NASA through the NASA Hubble Fellowship grant HST-HF2-51504.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., for NASA, under contract NAS5-26555. A. T. acknowledges support by NSF AST-2009884, AST-2107839, AST-1815304, AST-1911080, OAC-2031997, AST-2206471 and NASA 80NSSC22K0031 and 80NSSC21K1746 grants. K. D. A. acknowledges support provided by the NSF through award AST-2307668.